# A Multi-Level Superoptimizer for Tensor Programs

### v2

@@ 5. Probabilistic Equivalence Verifier
Mirage’s probabilistic equivalence verifier checks if a candidate $\mu$Graph is equivalent to the desired Lax subprogram. The key idea is to evaluate both on random inputs in two finite fields. Using finite fields instead of floating point numbers not only avoids floating point errors, but also leads to a strong theoretical guarantee: the probability of accepting a non-equivalent $\mu$Graph can be made arbitrarily low.
@@
For general programs, random tests can hardly provide any correctness guarantees. However, we show that for Lax programs (formally defined below), random testing provides a probabilistic correctness guarantee, and repeated tests can reduce the error probability to an arbitrarily small threshold.
@@
Prior work (wang2021pet) has applied a similar technique to check equivalence between tensor programs that only contain linear operators (e.g., matrix multiplication, convolution). We develop a random testing technique that also supports division and exponentiation, which are needed for many DNN optimizations (e.g., the RMSNorm example in §3).
@@
Mirage verifies equivalence between Lax $\mu$Graphs (linear, division, and an exponential) defined below. We introduce the main theoretical results in §5.1 and present Mirage’s verification methodology in §5.2.
Definition 1 (Lax $\mu$Graph).
@@
5.1. Theoretical Foundations
Without loss of generality, we assume a Lax $\mu$Graph $G$ takes $n$ input tensors and produces one output tensor. Our theoretical results can directly generalize to Lax $\mu$Graph with multiple outputs. Since each Lax $\mu$Graph includes linear operators, divisions, and at most one exponentiation along each path, the computation for each entry of the output tensor can be expressed in the following form (by using standard identities such as $\frac{\frac{a}{b}}{\frac{c}{d}}=\frac{ad}{bc}$, $\frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bd}$, $e^{x}e^{y}=e^{x+y}$):
(2) $\sum_{i=1}^{k}\frac{f_{i}}{g_{i}}e^{\left({\frac{h_{i}}{u_{i}}}\right)},$
where $f_{i}$, $g_{i}$, $h_{i}$, and $u_{i}$ ($1\leq i\leq k$) are polynomials over the entries of the input tensors.
The main theoretical result that underpins our randomized equivalence verification is the following theorem, which extends polynomial identity testing (PIT) (schwartz1980fast; zippel1979probabilistic) on finite fields to Lax $\mu$Graphs. Note that the difference of two Lax $\mu$Graphs is also of the form of Equation 2. Therefore, identity testing of two Lax $\mu$Graphs can be done by testing if an expression of that form is zero. Because of the use of exponentiation, we use two finite fields instead of one. [^5] We use two primes $p$ and $q$ for polynomial identity testing (schwartz1980fast; zippel1979probabilistic) outside and inside the exponents, respectively. The condition $q$ divides $p-1$ is to ensure the existence of $q$-th roots of unity in $\mathbb{Z}_{p}$.
@@ 5.1. Theoretical Foundations
Let $\mathbb{Z}_{p},\mathbb{Z}_{q}$ be the finite field of integers modulo $p$ and $q$, respectively, where $p,q$ are primes such that $q$ divides $p-1$. Let $d,k\in\mathbb{N}$ s.t. $dk^{4}<q$. Let $f_{1},\ldots,f_{k},g_{1},\ldots,g_{k}:\mathbb{Z}_{p}^{N}\to\mathbb{Z}_{p}$ be polynomials over $\mathbb{Z}_{p}$ and $h_{1},\ldots,h_{k},u_{1},\ldots,u_{k}:\mathbb{Z}_{q}^{M}\to\mathbb{Z}_{q}$ be polynomials over $\mathbb{Z}_{q}$, where the degrees of all polynomials are at most $d$. Let $\mathcal{G}$ be the set of $q$-th roots of unity in $\mathbb{Z}_{p}$. If $h_{i}/u_{i}\not\equiv h_{j}/u_{j}$ for all $i\neq j$, and $f_{i},g_{i},h_{i}\not\equiv 0$ for all $i$, then
$\Pr_{\mathbf{x}\leftarrow\mathbb{Z}_{p}^{N},\mathbf{y}\leftarrow\mathbb{Z}_{q}^{M},\omega\leftarrow\mathcal{G}}\left[\sum_{i=1}^{k}\frac{f_{i}(\mathbf{x})}{g_{i}(\mathbf{x})}\omega^{\frac{h_{i}(\mathbf{y})}{u_{i}(\mathbf{y})}}=0\wedge\mathcal{E}\right]\leq 1-\frac{1}{k}+o\left(\frac{1}{k}\right),$
where $\mathcal{E}$ is the event that $g_{i}(\mathbf{x}),u_{i}(\mathbf{y})\neq 0$ for all $i$.
5.2. Random Tests over Finite Fields
Table 3. Arithmetic operations for random testing. Mirage selects two prime numbers $p$ and $q$ such that $q$ divides $p-1$. $x_{p}$ and $x_{q}$ are values from $\mathbb{Z}_{p}$ and $\mathbb{Z}_{q}$, respectively.
@@ 5.2. Random Tests over Finite Fields
Add. $(x_{p},x_{q})$ $(y_{p},y_{q})$ $\big{(}x_{p}+y_{p}(\bmod{p}),x_{q}+y_{q}(\bmod{q})$)
Sub. $(x_{p},x_{q})$ $(y_{p},y_{q})$ $\big{(}x_{p}-y_{p}(\bmod{p}),x_{q}-y_{q}(\bmod{q})$)
Mul. $(x_{p},x_{q})$ $(y_{p},y_{q})$ $\big{(}x_{p}y_{p}(\bmod{p}),x_{q}y_{q}(\bmod{q})\big{)}$
Div. $(x_{p},x_{q})$ $(y_{p},y_{q})$ $\big{(}x_{p}y_{p}^{-1}(\bmod{p}),x_{q}y_{q}^{-1}(\bmod{q})\big{)}$
Exp. $(x_{p},x_{q})$ $\big{(}\omega^{x_{q}}(\bmod{p}),-\big{)}$
Mirage leverages Theorem 2 to probabilistically verify the equivalence of two $\mu$Graphs by performing random testing over the finite fields $\mathbb{Z}_{p}$ and $\mathbb{Z}_{q}$ as defined in Theorem 2. To check the equivalence of two $\mu$Graphs, Mirage first generates the input tensors, where each entry is uniformly sampled from $\mathbb{Z}_{p}\times\mathbb{Z}_{q}$. Mirage also samples $\omega$, which is used for exponentiation, uniformly from all the $q$-roots of unity in $\mathbb{Z}_{p}$. Mirage then evaluates the two $\mu$Graphs over the finite fields for the same inputs using the operations defined in Table 3. As explained in §5.1, $\mathbb{Z}_{p}$ and $\mathbb{Z}_{q}$ are used outside and inside the exponent, respectively. All operations except exponentiation are implemented via modular arithmetic on $\mathbb{Z}_{p}$ and $\mathbb{Z}_{q}$ separately. Exponentiation uses the value $x_{q}$ from $\mathbb{Z}_{q}$ and transforms it to a value in $\mathbb{Z}_{p}$ by computing $\omega^{x_{q}}(\bmod p)$. Note that in a Lax $\mu$Graph, exponentiation will be performed at most once along each path. Finally, Mirage checks whether the two $\mu$Graphs produce identical outputs. This process is repeated multiple times, and the two $\mu$Graphs pass the equivalence test if they pass all random tests. The following theorem, which follows from Theorem 2, shows that this process can yield an arbitrarily low error rate.
@@
Equivalent $\mu$Graphs always pass $\mu$Graph verification. For two non-equivalent $\mu$Graphs and a given probability threshold $0<\delta\leq 1$, the $\mu$Graphs pass all $\Omega(k\cdot\ln\frac{1}{\delta})$ random tests with probability at most $\delta$.
6. $\mu$Graph Optimizer
For each verified $\mu$Graph, Mirage’s $\mu$Graph optimizer maximizes its performance by further performing layout optimizations, operator scheduling, and memory planning, as shown in Figure 1. Mirage defers these $\mu$Graph optimizations after verification for two reasons. First, these optimizations do not affect the correctness of the generated $\mu$Graphs; omitting them when generating $\mu$Graphs reduces the search space Mirage must consider, since $\mu$Graphs with the same graph topology but different choices of tensor layouts, operator orders, or memory allocation plans are considered identical by the $\mu$Graph generator. Second, considering these optimizations after verification also reduces the search space for these optimizations, since the $\mu$Graph optimizer only needs to optimize $\mu$Graphs that are functionally equivalent to the input.
@@ 6. $\mu$Graph Optimizer
The $\mu$Graph optimizer considers possible data layouts for all intermediate tensors at the kernel, block, and thread levels and chooses the best combinations to maximize performance. We formulate layout selection as a constrained optimization problem and use an integer linear programming (ILP) algorithm to solve it optimally. Specifically, for each tensor $t$ and each possible layout $l$ for $t$, we introduce a boolean variable $B_{t,l}$ to indicate whether tensor $t$ uses layout $l$. Operators at the kernel, block, and thread levels may impose various constraints on tensor layouts. For example, to use kernels in the cuBLAS library (cublas) to perform matrix multiplication, the innermost dimension of the two input tensors must be among the last two dimensions. These restrictions are converted into a series of linear constraints on $B_{t,l}$. Different tensor layouts may lead to varying performance. For example, some input tensor layouts support bulk copies from device to shared memory, while others do not. Mirage introduces a cost function to model the performance of each operator under different layout choices. Mirage uses an off-the-shelf ILP solver (i.e., Z3 (z3)) to find an optimal layout strategy that satisfies all layout constraints while minimizing cost.
@@
In a $\mu$Graph, there are multiple topological orders to execute operators, and different orders may yield different performance. For a given input $\mu$Graph, the $\mu$Graph optimizer identifies an efficient strategy to schedule operators by minimizing thread-level synchronization within each thread block (i.e., __syncthreads() in CUDA). To achieve this goal, Mirage labels each node with a depth, defined as the length of the longest path from any input operator to that node. Mirage uses a dynamic programming algorithm to compute the depth of each node and schedules all operators in ascending order of their depths. This approach minimizes the number of thread-level synchronizations required in the generated CUDA kernel, as Mirage only needs to insert synchronization points between operators of different depths.
@@
A third class of post-verification optimizations is memory planning, which determines the memory offsets for all intermediate tensors at the kernel, block, and thread levels. Mirage formulates memory planning as a dynamic storage allocation task and exhaustively enumerates all allocation plans to discover an optimal strategy.
7. Implementation
Table 4. DNN benchmarks used in our evaluation.
@@ 7. Implementation
GQA Group-query attention LLaMA-3-70B (grattafiori2024llama3herdmodels)
QKNorm QK normalization with attention Chameleon-7B (chameleonteam2024chameleon)
RMSNorm RMS normalization with linear LLaMA-2-7B (llama2)
LoRA Low-rank adaptation GPT-3-7B-LoRA (llama-7b-lora)
GatedMLP Gated multi-layer perceptron Falcon-7B (falcon40b)
nTrans Normalized Tarnsformer nGPT-1B (loshchilov2024ngpt)
Mirage is implemented in 30K lines of code in C++, CUDA, and Python. Kernel operators are implemented with the cuDNN and cuBLAS libraries (cudnn; cublas), and block and thread operators are implemented using cuTLASS (cutlass) and CUDA functions. For each input tensor program, Mirage automatically generates and verifies potential $\mu$Graphs. For each verified $\mu$Graph, Mirage produces CUDA source code for all custom kernels of the $\mu$Graph and compiles the code into binary using the CUDA compiler. This approach enables just-in-time (JIT) compilation and deployment for general tensor programs, and the generated kernels can be directly integrated into a PyTorch program with a few lines of code changes. The SMT and ILP solvers of Mirage are implemented using Z3 4.12.6 (z3).
Our implementation supports the operators listed in Table 1. Mirage can be extended to include new operators, such as variants of convolution or matrix multiplication, at the kernel, block, and/or thread levels. To support a new linear operator, Mirage requires (1) a float-pointing implementation of the operator at the kernel, block, and/or thread levels, which is used by the $\mu$Graph optimizer to generate CUDA kernels, (2) an implementation of the operator over modular arithmetic (see §5), and (3) an extension to the abstract expression axioms $A_{\text{eq}}$ and $A_{\text{sub}}$ for it (see §4.3).
To utilize Theorems 2 and 3, random tests should be performed with sufficiently large prime numbers $p$ and $q$ and should be iterated multiple times. Our current implementation uses the largest values of $p$ and $q$ whose product fits in 16-bit integers (i.e., $p=227,q=113$) to perform the random testing on GPUs, leveraging Mirage’s GPU optimizations such as maintaining intermediate results in GPU shared memory, which allows Mirage to accelerate its search procedure. We also use a single random test without iterating it and compare all elements of the output tensors. We note that this equivalence verification procedure does not introduce false negatives. While false positives may be introduced, we have not observed any in practice. For these reasons, we consider this procedure to be sufficient for the search process and plan to add an additional verification step that provides the theoretical guarantees only for the best $\mu$Graph at the end of the optimization process.
@@
While Mirage can generate $\mu$Graphs for arbitrary input tensor programs, the probabilistic equivalence verifier is restricted to Lax programs and does not support certain DNN operators such as ReLU (relu). As an alternative, we have developed a solver-based verifier for arbitrary tensor programs. The verifier relies on user-provided mathematical properties of individual operators (e.g., linearity, associativity, commutativity, and distributivity) defined in first-order logic and uses these properties to verify equivalence using an automated theorem prover. Compared to the probabilistic equivalence verifier, the solver-based verifier supports more general programs, while requiring additional manual effort to specify the properties for each new operator. A detailed discussion of the solver-based verifier is beyond the scope of this paper.
8. Evaluation
8.1. Experimental Setup
Figure 7. Comparing Mirage with existing systems for 6 benchmarks on an A100 and an H100 GPU. The performance of all systems are normalized by Mirage (higher is better). Numbers above the Mirage bars show the speedup over the best baselines.
Since Mirage is a superoptimizer for Lax programs, we focus our evaluation on various DNN benchmarks commonly used by existing DNNs, each of which is a Lax program. These benchmarks provide the most fine-grained way to compare the performance of Mirage and existing systems. Table 4 shows the six benchmarks in our evaluation. GQA, RMSNorm, and GatedMLP are the main building blocks of large language models (LLMs). QKNorm introduces query-key normalization before attention to enhance model convergence (chameleonteam2024chameleon). LoRA enables low-rank adaptation for fine-tuning a DNN on different tasks. In addition, we also evaluate how Mirage-generated kernels can improve the end-to-end performance of DNNs, including Chameleon (chameleonteam2024chameleon), nGPT (loshchilov2024ngpt), LLaMA-3 (grattafiori2024llama3herdmodels), and LoRA (hu2021lora).
The experiments were performed on a NVIDIA 40GB A100 and a NVIDIA 40GB H100 GPU. All our benchmarks fit on a single GPU except GQA (used for LLaMA-2-70B), which is generally parallelized across four GPUs using tensor model parallelism (Megatron). Therefore, we evaluate GQA under the tensor model parallelism strategy (i.e., the 8 key-value heads are equally partitioned across four GPUs). Since the performance of Mirage and all baselines depends only on the shapes of the input tensors, we repeat each experiment 1,000 times using random inputs and report the average run time.

### v3

@@ 5 Probabilistic Equivalence Verifier
Mirage’s probabilistic equivalence verifier checks if a candidate $\mu$Graph is equivalent to the desired Lax program. The key idea is to evaluate both on random inputs in two finite fields. Using finite fields instead of floating point numbers not only avoids floating point errors but also provides a strong theoretical guarantee: the probability of accepting a non-equivalent $\mu$Graph can be made arbitrarily low.
@@
For general programs, random testing can hardly provide any correctness guarantee. However, we show that for Lax programs (formally defined below), random testing offers a probabilistic correctness guarantee, and repeated tests can reduce the error probability to an arbitrarily small threshold.
@@
Prior work [46] has applied a similar technique to check equivalence between tensor programs that contain only linear operators (e.g., matrix multiplication, convolution). We develop a random testing technique that also supports division and exponentiation, which are needed for many DNN optimizations (e.g., the RMSNorm example in §3).
@@
Mirage verifies equivalence between Lax $\mu$Graphs (linear, division, and an exponentiation) defined below. We introduce the main theoretical results in §5.1 and present Mirage’s verification methodology in §5.2.
Definition 5.1 (Lax $\mu$Graph).
@@
5.1 Theoretical Foundations
Without loss of generality, we assume a Lax $\mu$Graph $G$ takes $n$ input tensors and produces one output tensor. Our theoretical results directly generalize to Lax $\mu$Graph with multiple outputs. Since each Lax $\mu$Graph includes linear operators, divisions, and at most one exponentiation along each path, the computation for each entry of the output tensor can be expressed in the following form (by using standard identities such as $\frac{\frac{a}{b}}{\frac{c}{d}}=\frac{ad}{bc}$, $\frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bd}$, $e^{x}e^{y}=e^{x+y}$):
$\frac{\sum_{i=1}^{k}f_{i}\exp(g_{i}/h_{i})}{\sum_{i=1}^{k^{\prime}}f^{\prime}_{i}\exp(g^{\prime}_{i}/h^{\prime}_{i})}$ (2)
where $f_{i}$, $g_{i}$, $h_{i}$, $f_{j}^{\prime}$, $g_{j}^{\prime}$ and $h_{j}^{\prime}$ ($1\leq i\leq k$, $1\leq j\leq k^{\prime}$) are polynomials over the entries of the input tensors.
The main theoretical result that underpins our randomized equivalence verification is the following theorem, which extends polynomial identity testing (PIT) [37, 54] on finite fields to Lax $\mu$Graphs. Note that the difference of two Lax $\mu$Graphs is also of the form of Equation 2. Therefore, identity testing of two Lax $\mu$Graphs reduces to testing if an expression of that form is zero. Due to the presence of exponentiation, we use two finite fields instead of one. [^5] We use two primes $p$ and $q$ for polynomial identity testing [37, 54] outside and inside the exponents, respectively. The condition $q$ divides $p-1$ is to ensure the existence of $q$-th roots of unity in $\mathbb{Z}_{p}$.
@@ 5.1 Theoretical Foundations
Let $P$ be a function of the form described in Equation 2, where $f_{i},g_{i},h_{i},f_{i}^{\prime},g_{i}^{\prime},h_{i}^{\prime}$ are non-zero polynomials of degree at most $d$ with integer coefficients between $[-w,w]$. Let $p,q$ be primes such that $q\mid p-1$ and $q>2w$. Let $\mathcal{G}$ be the set of $q$-th roots of unity in $\mathbb{Z}_{p}$. If $P$ is not a zero function, then [27]
$\Pr_{(\vec{u},\vec{v},\omega)\leftarrow\mathbb{Z}_{p}^{N}\times\mathbb{Z}_{q}^{N}\times\mathcal{G}}\left[\frac{\sum_{i=1}^{k}f_{i}(\vec{u})\omega^{g_{i}(\vec{v})/h_{i}(\vec{v})}}{\sum_{i=1}^{k^{\prime}}f^{\prime}_{i}(\vec{u})\omega^{g^{\prime}_{i}(\vec{v})/h^{\prime}_{i}(\vec{v})}}\right]\leq 8dk^{4}/q+q^{-1/k^{2}}.$
5.2 Random Tests over Finite Fields
Table 3: Arithmetic operations for random testing. Mirage selects two prime numbers $p$ and $q$ such that $q$ divides $p-1$. $x_{p}$ and $x_{q}$ are values from the finite fields $\mathbb{Z}_{p}$ and $\mathbb{Z}_{q}$, respectively. The notation $x^{-1}$ and $\sqrt{x}$ represents the multiplicative inverse and square root of $x$ in the corresponding finite field. Specifically, $xx^{-1}\bmod p=1$ and $\sqrt{x}\sqrt{x}\bmod p=x$.
@@ 5.2 Random Tests over Finite Fields
Add. $(x_{p},x_{q})$ $(y_{p},y_{q})$ $\big{(}(x_{p}+y_{p})\bmod{p},(x_{q}+y_{q})\bmod{q}$)
Sub. $(x_{p},x_{q})$ $(y_{p},y_{q})$ $\big{(}(x_{p}-y_{p})\bmod{p},(x_{q}-y_{q})\bmod{q}$)
Mul. $(x_{p},x_{q})$ $(y_{p},y_{q})$ $\big{(}x_{p}y_{p}\bmod{p},x_{q}y_{q}\bmod{q}\big{)}$
Div. $(x_{p},x_{q})$ $(y_{p},y_{q})$ $\big{(}x_{p}y_{p}^{-1}\bmod{p},x_{q}y_{q}^{-1}\bmod{q}\big{)}$
Exp. $(x_{p},x_{q})$- $\big{(}\omega^{x_{q}}\bmod{p},-\big{)}$
Sqrt. $(x_{p},x_{q})$- $(\sqrt{x_{p}},\sqrt{x_{q}})$
Mirage leverages Theorem 2 to probabilistically verify the equivalence of two $\mu$Graphs by performing random testing over the finite fields $\mathbb{Z}_{p}$ and $\mathbb{Z}_{q}$ as defined in Theorem 2. To check the equivalence of two $\mu$Graphs, Mirage first generates input tensors, with each entry uniformly sampled from $\mathbb{Z}_{p}\times\mathbb{Z}_{q}$. Mirage also samples $\omega$ uniformly from the set of $q$-roots of unity in $\mathbb{Z}_{p}$, which is used for exponentiation. Mirage then evaluates the two $\mu$Graphs on these inputs using the operations defined in Table 3. As explained in §5.1, $\mathbb{Z}_{p}$ and $\mathbb{Z}_{q}$ are used for computations outside and inside the exponent, respectively. All operations except exponentiation are implemented via modular arithmetic in $\mathbb{Z}_{p}$ and $\mathbb{Z}_{q}$ independently. For exponentiation, Mirage uses the value $x_{q}$ from $\mathbb{Z}_{q}$ and computes $\omega^{x_{q}}\bmod p$ to obtain a result in $\mathbb{Z}_{p}$.
Note that in a Lax $\mu$Graph, exponentiation is performed at most once along each path. Finally, Mirage checks whether the two $\mu$Graphs produce identical outputs. This process is repeated multiple times, and the two $\mu$Graphs are considered equivalent if they pass all random tests. The following theorem, which follows from Theorem 2, shows that this process can yield an arbitrarily low error rate.
@@
Equivalent $\mu$Graphs always pass $\mu$Graph verification. For two non-equivalent $\mu$Graphs and a given probability threshold $0<\delta\leq 1$, the $\mu$Graphs pass all $\Omega(\frac{k^{2}}{\ln q}\cdot\ln\frac{1}{\delta})$ random tests with probability at most $\delta$.
Numerical stability.
While the theorem bridges finite fields and real-number computations, discrepancies can arise between real-number computations and floating-point operations, particularly involving overflow or underflow due to large intermediate values. Mirage employs floating-point tests to filter out $\mu$Graphs with significant numerical errors.
6 $\mu$Graph Optimizer
For each verified $\mu$Graph, Mirage’s $\mu$Graph optimizer maximizes its performance by further performing layout optimization, operator scheduling, and memory planning, as shown in Figure 1. Mirage defers these $\mu$Graph optimizations until after verification for two reasons. First, these optimizations do not affect the correctness of the generated $\mu$Graphs; omitting them when generating $\mu$Graphs reduces the search space Mirage must consider, since $\mu$Graphs with the same graph topology but different choices of tensor layouts, operator orders, or memory allocation plans are considered identical by the $\mu$Graph generator. Second, applying these optimizations after verification also reduces the search space for these optimizations, since the $\mu$Graph optimizer only needs to optimize $\mu$Graphs that are functionally equivalent to the input.
@@ 6 $\mu$Graph Optimizer
The $\mu$Graph optimizer explores possible data layouts for all intermediate tensors at the kernel, block, and thread levels and chooses the best combinations to maximize performance. We formulate layout selection as a constrained optimization problem and solve it optimally using an integer linear programming (ILP) algorithm. Specifically, for each tensor $t$ and each possible layout $l$ for $t$, we introduce a boolean variable $B_{t,l}$ to indicate whether tensor $t$ uses layout $l$. Operators at the kernel, block, and thread levels may impose various constraints on tensor layouts. For example, to use kernels from the cuBLAS library [16] for matrix multiplication, the innermost dimension of the two input tensors must be among the last two dimensions. These restrictions are converted into a series of linear constraints on $B_{t,l}$. Different tensor layouts may lead to varying performance. For example, some input tensor layouts support bulk copies from device to shared memory, while others do not. Mirage introduces a cost function to model the performance of each operator under different layout choices. Mirage uses an off-the-shelf ILP solver (i.e., Z3 [18]) to find an optimal layout strategy that satisfies all layout constraints while minimizing cost.
@@
In a $\mu$Graph, there are multiple topological orders to execute operators, and different orders may yield different performance. For a given input $\mu$Graph, the $\mu$Graph optimizer identifies an efficient strategy to schedule operators by minimizing thread-level synchronization within each thread block (i.e., __syncthreads() in CUDA). To achieve this goal, Mirage labels each node with a depth, defined as the length of the longest path from any input operator to that node. Mirage uses a dynamic programming algorithm to compute the depth of each node and schedules all operators in ascending order of their depths. This approach minimizes the number of thread-level synchronizations required in the generated CUDA kernel, as Mirage only needs to insert synchronization points between operators with different depths.
@@
A third class of post-verification optimizations is memory planning, which determines memory offsets for all intermediate tensors at the kernel, block, and thread levels. Mirage formulates memory planning as a dynamic storage allocation problem and exhaustively enumerates all possible allocation plans to discover an optimal strategy.
7 Implementation
Table 4: DNN benchmarks used in our evaluation.
@@ 7 Implementation
GQA Group-query attention LLaMA-3-70B [41]
QKNorm QK normalization with attention Chameleon-7B [40]
RMSNorm RMS normalization with linear LLaMA-2-7B [44]
LoRA Low-rank adaptation GPT-3-7B-LoRA [6]
GatedMLP Gated multi-layer perceptron Falcon-7B [10]
nTrans Normalized Tarnsformer nGPT-1B [28]
Mirage is implemented in 30K lines of code in C++, CUDA, and Python. Kernel operators are implemented with the cuDNN and cuBLAS libraries [15, 16], and block and thread operators are implemented using cuTLASS [2] and CUDA PTX. For each input tensor program, Mirage automatically generates and verifies potential $\mu$Graphs. For each verified $\mu$Graph, Mirage produces CUDA source code for all custom kernels of the $\mu$Graph and compiles the code into binary using the CUDA compiler. This approach enables just-in-time (JIT) compilation and deployment for general tensor programs, and the generated kernels can be directly integrated into a PyTorch program with a few lines of code changes. Mirage’s SMT and ILP solvers are implemented using Z3 4.12.6 [18].
Our implementation supports the operators listed in Table 1. Mirage can be extended to include new operators, such as variants of convolution or matrix multiplication, at the kernel, block, and/or thread levels. To support a new linear operator, Mirage requires (1) a float-pointing implementation of the operator at the kernel, block, and/or thread levels, which is used by the $\mu$Graph optimizer to generate CUDA kernels; (2) an implementation of the operator over modular arithmetic (see §5); and (3) an extension to the abstract expression axioms $A_{\text{eq}}$ and $A_{\text{sub}}$ for the operator (see §4.3).
To utilize Theorems 2 and 3, random tests should be performed with sufficiently large prime numbers $p$ and $q$ and iterated multiple times. Our current implementation uses the largest values of $p$ and $q$ whose product fits in 16-bit integers (i.e., $p=227,q=113$) to run these random tests on GPUs. We leverage Mirage’s GPU optimizations–such as keeping intermediate results in shared memory–to accelerate the search procedure. We also perform a single random test without iterating it and compare all elements of the output tensors. We note that this equivalence verification procedure does not introduce false negatives. While it could, in theory, introduce false positives, we have not observed any in practice. For these reasons, we consider this procedure sufficient for the search process and plan to add a final verification step that provides the theoretical guarantees only for the best $\mu$Graph at the end of the optimization process.
@@
While Mirage can generate $\mu$Graphs for arbitrary tensor programs, the probabilistic equivalence verifier is limited to Lax programs and does not support certain DNN operators such as ReLU [32]. As an alternative, we have developed a solver-based verifier for arbitrary tensor programs. The verifier relies on user-provided mathematical properties of individual operators (e.g., linearity, associativity, commutativity, and distributivity) defined in first-order logic and uses these properties to verify equivalence using an automated theorem prover. Compared to the probabilistic equivalence verifier, the solver-based verifier supports more general programs, while requiring additional manual effort to specify the properties of each new operator. A detailed discussion of the solver-based verifier is beyond the scope of this paper.
8 Evaluation
8.1 Experimental Setup
Figure 7: Comparing Mirage with existing systems for 6 benchmarks on an A100 and an H100 GPU. The performance of all systems are normalized by Mirage (higher is better). Numbers above the Mirage bars show the speedup over the best baselines.
Since Mirage is a superoptimizer for Lax programs, we focus our evaluation on various DNN benchmarks commonly used in existing DNNs, each of which is a Lax program. These benchmarks provide the most fine-grained way to compare the performance of Mirage and existing systems. Table 4 shows the six benchmarks in our evaluation. GQA, RMSNorm, and GatedMLP are the main building blocks of large language models (LLMs). QKNorm introduces query-key normalization before attention to enhance model convergence [40]. LoRA enables low-rank adaptation for fine-tuning a DNN on different tasks. We use a context length of 8K for GQA and 4K for QKNorm, corresponding to the maximum supported by LLaMA-3-70B [41] and Chameleon-7B [40], respectively. In addition, we also evaluate how Mirage-generated kernels improve the end-to-end performance of full DNNs, including Chameleon [40], nGPT [28], LLaMA-3 [41], and LoRA [22].
The experiments were conducted on NVIDIA A100 and H100 GPUs, each with 40GB of memory. All our benchmarks fit on a single GPU except GQA (used for LLaMA-2-70B), which is generally parallelized across four GPUs using tensor model parallelism [39]. Therefore, we evaluate GQA under this parallelism strategy, where the eight key-value heads are equally partitioned across four GPUs. Since the performance of Mirage and all baselines depends only on the shapes of the input tensors, we repeat each experiment 1,000 times using random inputs and report the average run time.